OpenStudy (anonymous):

find x intercepts of 4x^2-8x+4

OpenStudy (campbell_st):

let your quadratic equal zero and solve. \[4x^2 - 8x + 4 = 0\] you can take out a factor of 4 \[4(x^2 - 2x + 1) = 0\] just factor the quadratic to then solve for x to find the x-intercept

OpenStudy (anonymous):

I'm getting that there are two x intercepts correct?

OpenStudy (campbell_st):

well yes there are but you'll find its a repeated value. you are factoring a perfect square

OpenStudy (anonymous):

Ok, I am a little confused. Can you show me by working it out in steps, maybe then I'll see it and understand it.

OpenStudy (campbell_st):

ok, what are the factors of 1 that add to -2, they are both negative.

OpenStudy (anonymous):

1+-3

OpenStudy (campbell_st):

if you multiply them you get - 3 its -1 * -1 = 1 -1 + -1 = -2 does that make sense?

OpenStudy (anonymous):

Yes that makes more sense, thank you.

OpenStudy (anonymous):

I was confused when the answers did not offer a repeated value.

OpenStudy (campbell_st):

well I think you are looking at \[4(x -1)^2 = 0\] what value of x makes the equation true, that is the x- intercept.

OpenStudy (anonymous):

ah...ok